OpenStudy (anonymous):

im having a really hard time solving and graphing liner equations... plz help!!

OpenStudy (anonymous):

??..wat you don't understand?

OpenStudy (amistre64):

wait til you have to make a curve :)

OpenStudy (anonymous):

true

OpenStudy (anonymous):

i dnt know how to solve them

OpenStudy (anonymous):

do u mean finding the y and x intercepts?

OpenStudy (amistre64):

give us something to work with....

OpenStudy (anonymous):

like y + x = -7

OpenStudy (amistre64):

Do you know all the parts to this equation: y = mx + b

OpenStudy (anonymous):

no i dont understand that part

OpenStudy (amistre64):

y stands for the value of your "y" coordinate. m stands for the way the line slopes; called the slope. x stands for the value of the "x" coordinate. b is the value of where the line crosses the y-axis. does any of that make sense?

OpenStudy (anonymous):

it makes some sence... i just get confused when it comes to solving more then graphing.

OpenStudy (amistre64):

as long as you can get an equation to resemble this construction (y = mx+b) then you can solve for any point on the line. Lets take your equation you gave us for an example...

OpenStudy (amistre64):

y + x = -7 ; do you know how to make this look more like (y=mx+b)?

OpenStudy (anonymous):

do u have yahoo messanger??

OpenStudy (amistre64):

nope...

OpenStudy (anonymous):

ok sorry my daughter is messinga round

OpenStudy (anonymous):

so when i read in the book its so confusing on how to solve these probelms can you show me step by step what to do?

OpenStudy (amistre64):

I can try: In essense; we want to know what values to make "y" for any given value of "x". In your equation; we have: y + x = -7 use algebraic techniques to solve for "y"; which in this case just means subtracting (x) from both sides to get: y +x-x = -7-x y = -x -7 ; which resembles our (y = mx + b) Do you understand it so far?

OpenStudy (amistre64):

Once we have any two points that we can connect together; we can form a straight line. Many times, those two points will be where the line crosses the y-axis; and where the line crosses the x-axis. These crossing points are called "intercepts" becasue it is where the line actually gets intercepted by the x or y axis.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

What two points (called coordinates) can you come up with using : y = -x -7 in order to connect and make a line?